OpenStudy (anonymous):

3sqrt{-27} is the answer 9??

OpenStudy (anonymous):

\[3\sqrt{-27}\]

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

sqrt(-27) = 3(i)sqrt(3) The answer is 9 (i) (sqrt(3))

OpenStudy (anonymous):

i think my answers are wrong lol i cant multiply lol

OpenStudy (anonymous):

i is imaginary number, which is = to sqrt(-1)

OpenStudy (anonymous):

\[9\sqrt{3}\] would it look like that as an answer?

OpenStudy (anonymous):

\[3i \sqrt{3}\]

OpenStudy (anonymous):

sorry 9i no 3i

OpenStudy (anonymous):

ive never seen an i in the answer thats a new one on me

OpenStudy (anonymous):

i is an imaginary number you cant take the sqrt of a negative number its not possible unless you use imaginary number (i)

OpenStudy (anonymous):

so do i have to put the i in my answer? i guess thats where im confused is how to write it as an answer

OpenStudy (anonymous):

yes it should be \[9i \sqrt{3}\] with out the i there cant be an answer, now in class if you haven't learned (i) yet then i'm not sure what you are suppose to put down but that is what i would write as my answer

OpenStudy (anonymous):

hope that helps

OpenStudy (anonymous):

i am going to make a bet that the question is actually \[\sqrt[3]{-27}\] in which case the answer is \[-3\]

OpenStudy (anonymous):

yea we havent learned anything about an i thats why im lost on that

OpenStudy (anonymous):

@tena this is not "three square root of minus 27" it is "cubed root of -27"

OpenStudy (anonymous):

i had -3 as an answer earlier lol

OpenStudy (anonymous):

not the "square root" but the "cubed root" and it is -3 because \[(-3)^3=-27\]

OpenStudy (anonymous):

ohhhh....well that makes a lot more sense if its the cube root of (-27) sense you have not learned the number (i) yet

OpenStudy (anonymous):

yea im sorry it is. I typed it in wronged but worked it out right lol im so sorry

OpenStudy (anonymous):

no problem sorry i confused you with imaginary numbers